Steve, 18 years old, is three times as old as his brother Daniel. How old will Steve be when he will be twice as old as Daniel?
Answer by ewatrrr(24785)

Steve 18, Daniel 6
x number of years before Steve will be twice as old as Daniel
18+x = 2(6+x)
6 years = x, number of years hence. Then, Steve will be 24.
And...checking
24 = 2*12